OpenStudy (anonymous):

Please help me with this problem. I 5a+3 I=5 A.) 1,2 B.)2/5,-8/5 C.)2/5,8/5 D.)1,-1

OpenStudy (anonymous):

sorry please forget my previous comment. i would answer b

OpenStudy (anonymous):

lol its okay... ithought it was c but i am going to put B thanks :)

OpenStudy (anonymous):

based upon abs(5a+3) = 5a+3 if 5a+3>=0 and abs(5a+3) = -5a-3 if 5a+3 <=0

OpenStudy (anonymous):

okay so yes it would be b Can i ask you another one?

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

I 6c+5 I=4

OpenStudy (anonymous):

in this case first let 6c+5=4 for LHS >=0 and we get c=-1/6

OpenStudy (anonymous):

Chocies A.)-1/6,3/2 B.)-1/6,-3/2 C.)2/3,4/3 D.) 2/3,2/3

OpenStudy (anonymous):

then for the case LHS <0 we get -6c-5=4 which gives c=-3/2

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

can you see the method now?

OpenStudy (anonymous):

yes thanks so much :)